Circular No. 6099

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

HD 5980 R. Barba and V. Niemela, Observatório Astronómico, La Plata, com-

municate: “Spectral observations (360–700 nm) obtained between Oct. 24 and 27 with the REOSC échelle spectrograph + CCD attached to the 2.15m telescope at CASLEO Observatory, San Juan, show that the brightest Wolf-Rayet eclipsing binary star in the SMC, HD 5980 (α = 0^{h}58^{m}.62, δ = −72^{o}17′.9, equinox 1975.0; Azzopardi and Vigneau 1982, A.Ap.Suppl. 50, 291), has brightened considerably and completely changed its spectrum, which is now similar to η Car-type variables, with strong P-Cyg lines of He I and H Balmer, as well as Fe III and N II lines. From flux-calibrated spectra we estimate that the continuum V brightened by ∼ 1 mag with respect to previous observations obtained with the same telescope in Jan. and June, when it was at V = 11.0, already 0.8 mag brighter than published values (cf. Azzopardi and Vigneau 1982, ibid.). HD 5980 appears to be currently undergoing an η Car-type outburst.”

EXO 1846−031

H. Inoue, F. Nagase, M. Ishida and T. Sonobe, on behalf of the ASCA team, communicate: “Following the report by Zhang et al. (IAUC 6096), ASCA observed EXO 1846−031 between Oct. 22.326 and 22.785 UT with an exposure time of ∼ 12 000 s. No bright x-ray source was seen in the ASCA GIS field within 20′ of α = 18^{h}49^{m}.1, δ = −3^{o}08′ (equinox 2000.0). The 5σ upper limit of the 2–10-keV flux at the position of EXO 1846−031 (Parmar et al. 1993, A.Ap. 279, 179) is 0.01 mCrab (2×10^{−13} erg cm^{−2} s^{−1}, assuming the Crab Nebula spectrum). The upper limit increases to 0.04 mCrab at the rim of the GIS field. This suggests either that the outburst of EXO 1846−031 decayed very quickly or that the outburst detected by BATSE was not from this source.”

PERIODIC COMET TEMPEL 1 (1993c) Total visual magnitude estimates: Sept. 3.16 UT, 11.5 (C. S. Morris,

Lockwood Valley, CA, 0.26-m reflector); 4.16, 11.5 (Morris, Pine Mountain Club, CA, 0.26-m reflector).

SATURN

Corrigendum. On IAUC 6098, for f, 13.800 read f, 23.800.

1994 November 2 Gareth V. Williams